### v2

# Causal Models in Requirement Specifications for Machine Learning: A vision

Abstract.
Specifying data requirements for machine learning (ML) software systems remains a challenge in requirements engineering (RE). This vision paper explores causal modelling as an RE activity that allows the systematic integration of prior domain knowledge into the design of ML software systems. We propose a workflow to elicit low-level model and data requirements from high-level prior knowledge using causal models. The approach is demonstrated on an industrial fault detection system. This paper outlines future research needed to establish causal modelling as an RE practice.
AI Engineering, Causal Modelling, Data Requirements, Requirements Engineering
copyright: ccccs: Software and its engineering Requirements analysisccs: Software and its engineering Reusabilityccs: Software and its engineering Software prototyping
1. Introduction
Rahimi et al. called for more attention towards the ability of specifying software with machine learning (ML) components (Rahimi et al., 2019). Many industrial applications require robustness of ML models against changes in input data distribution (Borg et al., 2019). A key reason for lacking robustness is the difficulty of specifying ML models, because “if input and/or output data are high-dimensional, both defining preconditions and detailed function specifications are difficult”(Kuwajima et al., 2020). Robustness against context changes can only be tested if the expected operational context is explicitly defined, for instance through contextual requirements(Knauss et al., 2014, 2016). However, assumptions about the operational context are often implicit in the design process (Mitchell et al., 2021), such as in the selection of the training dataset. Recent surveys on requirements engineering (RE) confirm that specifying training data for ML models remains an open challenge (Pei et al., 2022; Ahmad et al., 2023; Franch et al., 2023). Current RE techniques struggle to translate high-level functional and non-functional requirements into data requirements (Alves et al., 2023; Villamizar et al., 2024). This leads to an underspecification causing variability in implementation choices and a lack of robustness against context changes (Fantechi et al., 2018).
A possible way to address underspecification is reasoning about expected causal relationships in the ML system’s operational context. Typically, ML cannot infer causality from data alone (Pearl, 2019). An ML model learns a probabilistic representation from data that seems compatible in a training context, but its performance might deviate drastically in a different operational context as statistical correlations do not capture true causal mechanisms (D’Amour et al., 2022). Addressing this challenge requires incorporating prior domain knowledge and causal reasoning into the design of ML systems.
This vision paper proposes causal modelling to communicate prior knowledge about causal relations in the operational context. We argue that by formulating prior domain knowledge as causal models we can derive requirements towards data, as well as deduce rules for runtime verification. This will lead to causally motivated requirements specifications for software with ML.
Objective of this vision paper
First, we outline our vision of integrating causal modelling as an RE activity for ML systems. Then, we illustrate its application in eliciting data requirements for an industrial prototype of an ML-based cooling fault-detection system for electric motors. Finally, we discuss a research agenda to explore the potential of causal modelling as an RE activity for ML systems.
2. Related Work
The potential of using causal modelling as part of RE activities is not yet fully explored (Giamattei et al., 2024). Fischbach et al. proposed an NLP-based process to extract and structure causal relationships from natural language (Fischbach et al., 2020, 2021). A tree recursive neural network (TRNN) model was trained to detect causality in natural language requirements using logical markers such as conjunctions and negations (Jadallah et al., 2021). They further developed an approach to converts extracted causal relationships into a DAG-like structure to automatically generate test cases (Fischbach et al., 2023). Maier et al. proposed modelling cause-effect relationships as part of scenario-based testing for automotive system safety (Maier et al., 2022). Maier et al. also introduced the concept of “CausalOps”, an industrial lifecycle framework for causal models (Maier et al., 2024). Gren et Brentsson Svensson proposed Bayesian Data Analysis (BDA) to evaluate the outcome of experiments on the effect of obsolete requirements on software effort estimation (Gren and Berntsson Svensson, 2021). Similarly, Frattini et al. investigated the impact of requirements quality defects on domain modelling by using BDA and causal reasoning in a in a controlled experiment (Frattini et al., 2025). While the latter two studies do not use causal modelling as an explicit RE activity, these studies demonstrates the potential of applying causal reasoning to RE activities.
3. Causal modelling as an RE activity
In a typical ML development pipeline, causal modelling would be a step between problem definition and data collection as it allows to formalise domain knowledge, identify relevant variables, and refine data requirements by distinguishing causal relationships from mere correlations before collecting the training data. Particularly, graphical causal models in the form of directed acyclic graph (DAG) allow to communicate explicitly assumed directions of causality and assumptions about confounders, i.e., situations in which a variable $Z$ is associated to two random variables $X_{1}$ and $X_{2}$ such that a spurious relationship between $X_{1}$ and $X_{2}$ can be observed: $X_{1}\,\leftarrow\,Z\,\rightarrow\,X_{2}$.
Figure 1. A proposed workflow for Causal RE
Figure 1 outlines a proposed workflow. The workflow bases on the principle of causal factorisation (Schölkopf et al., 2021):
(1) $p(X_{1},\ldots,X_{n})=\prod_{i=1}^{n}p\left(X_{i}|\mathbf{PA}_{i}\right),$
where $\mathbf{PA}_{i}$ denotes the set of parents (variables that have a direct causal effect) of a variable $X_{i}$ in the DAG. Causal factorisation implies that an observed joint distribution of interest can be decomposed into a product of conditional distributions, where each term corresponds to a causal mechanism.
Step 1) Identify individual causal mechanisms: The aim is to pinpoint specific cause-effect pathways informed based on high-level requirements, prior domain knowledge, and context assumptions.
Step 2) Update causal graph: Once a causal mechanism is identified, the relevant observable and latent variables are determined, and a causal graph is updated to include these variables along with the assumed directions of cause-and-effect relationships.
Step 3) Perform d-separation and extract requirements: With the causal model, d-separation [^1] Due to space constraints, background on d-separation is omitted but can be found in (Pearl, 2019). allows to identify variables that are needed to block “non-causal” association paths. Taking the example from above, in $X_{1}\leftarrow Z\rightarrow X_{2}$, there is a “non-causal” path between $X_{1}$ and $X_{2}$. If the ML model can condition on $Z$ (assuming $Z$ is observable), $X_{1}$ and $X_{2}$ become d-separated, closing the “non-causal” path. This is an example of a resulting data requirement: $Z$ must be included in the training dataset to avoid learning a spurious correlation between $X_{1}$ and $X_{2}$. Additionally, Step 3 provides independence criteria based on global Markov properties: If $X_{1}$ and $X_{2}$ are d-separated by $Z$, they are conditionally independent given $Z$, i.e., $X_{1}\perp X_{2}\,|\,Z$. This provides testable criteria to verify prior knowledge and assumptions encoded in the causal graph.
Step 4) Check consistency and observability: The graphical causal model must be checked for cyclic dependencies because a variable cannot be its own cause (Glymour et al., 2016). Furthermore, variables needed to block “non-causal” paths must be observable. If this is not the case, the system must be adjusted to enable their observation or suitable instrument variables must be identified (Angrist et al., 1996).
The resulting causal graph becomes part of an ML specification because it encodes the assumed causal structures, prior knowledge, and operational context, from which data and model requirements, as well as testing criteria, are derived.
4. Demonstration on industrial prototype
We demonstrate the use of causal modelling as an RE activity on an industrial prototype use case, specifically a system for detecting faults in the cooling system of electric motors.
Methodology:
We held three workshops with two Siemens engineers and two academic researchers to explore using causal models for requirements specification in the second half of 2022. The researchers introduced causal models with examples like temperature $\leftarrow$ sunrise $\rightarrow$ birds chirping and explaining key concepts such as confounding, colliders, and d-separation using for example the back-door criteria. The company experts then presented the prototype system, and prior knowledge rules were formalised together by identifying causal mechanisms and updating the causal model iteratively with each newly found causal mechanism. We then applied d-separation to close non-causal paths between the exposure (i.e., a cooling fault) and the outcome (i.e., the classification result) which resulted in data and model requirements to ensure the ML model controls for potential confounding.
Description of demonstration case:
The demonstration case, provided by Siemens, is a motor diagnostic application for monitoring electrical motors using an attachable sensor device. Initially, the system detected cooling faults from vibrations caused by mechanical faults, such as a broken fan blade. The new device will use an ML model to detect faults based on multiple sensor inputs. The high-level functional requirement is:
FR-1::
GIVEN indoor operational environment WHEN the cooling system is abnormal THEN an alarm should be raised.
The following prior knowledge of the company engineers was considered for identifying causal mechanisms:
PK-1::
A fault in the cooling system can affect the magnetic flux by changing the temperature of the rotor material and thus affecting the electrical resistance.
PK-2::
Mechanical faults of the fan can reduce the available airflow.
PK-3::
Mechanical faults cause vibrations of the system.
PK-4::
Environmental temperature has an influence on the temperature signal because the sensor is mounted outside the motor.
PK-5::
Unmeasured sensor disturbances exist in general.
Results:
The resulting causal model for the motor diagnostic use case is shown in Figure 2.
Figure 2. DAG for the motor diagnostic use case. Gray-background nodes are latent (unobservable) variables, while white-background nodes are observable at runtime.
Explanations for the variables and their relations to the prior knowledge are provided in Table 1.
Table 1. Variables for motor diagnostic use case.
Variable Related PKs Definition
Cooling Fault PK1, PK2 Fan system status
$Q$ PK2 Max. possible airflow
Mech. Fault PK2, PK3 Mechanical fault of motor
$P_{M}$ PK1 Mechanical power
$R_{1}$ PK1 Electrical (inner) losses
$T_{E}$ PK4 Environmental temperature
$U_{X}$ PK5 Unmeasured noises
$T$ ($T_{s}$) Surface temperature (measured)
$H$ ($H_{s}$) Magnetic Flux (measured)
$V$ ($V_{s}$) Vibrations (measured)
The causal graph in Figure 2 includes three causal mechanisms between Cooling Fault occurrence and Classification whether or not a cooling fault has occurred:
Temperature mechanism: A cooling fault increases the motor’s surface temperature $T$ (via the core temperature), measured by the temperature sensor $T_{s}$, which can be used to classify a cooling fault.
Magnetic flux mechanism: A cooling fault changes the inner resistance (via the core temperature), which affect the magnetic flux $H$. This is measured by the fluxmeter $H_{s}$ for classification.
Mechanical power mechanism: A cooling fault changes the magnetic flux $H$, which affects the mechanical power $P_{M}$ and surface temperature $T$. The latter is measured by the sensor $T_{s}$ for classification.
Two confounding paths were identified:
Mechanical fault confounding: A mechanical fan blade fault can reduce the available airflow $Q$ causing a cooling fault and vibrations $V$, which are measured by a vibration sensor $V_{s}$ for classification.
Environmental temperature confounding: A sudden change in environment temperature $T_{E}$ can temporarily limit cooling without indicating a fault and it affects the surface temperature $T$.
Data and model requirements:
We checked which variables must be observed and controlled for to close non-causal paths between cooling fault occurrence and classification of a cooling fault, which resulted in the requirements listed in Table 2. Vibration data $V_{s}$ alone is insufficient to detect cooling faults, as not all mechanical faults lead to a cooling fault (RQ-D1, RQ-M1). Instead, data on temperature and magnetic flux mechanisms should be included (RQ-M2). [^2] In fact, vibration data may be unnecessary for detecting cooling faults unless it is desired to distinguish mechanical from non-mechanical causes. An additional sensor should record the environmental temperature $T_{E}$ to control for confounding (RQ-D2). Sensor noise must also be represented in the training data (RQ-D3).
Table 2. Requirements derived from causal graph
ID Requirement (RQ-D: Data Req., RQ-M: Model Req.)
RQ-D1 Training data shall include cases where mechanical faults cause vibrations $V$ without leading to cooling faults .
RQ-D2 The occurrence of cooling faults shall be conditioned on different environmental temperatures $T_{E}$ such that the model can learn the confounding influence of $T_{E}$.
RQ-D3 Measurements shall include characteristic sensor noise.
RQ-M1 Cooling faults shall not be classified based on vibration data $V_{s}$ alone.
RQ-M2 The input layer shall accept temperature, magnetic flux, and vibration measurements.
Testing and runtime checks:
The causal graph in Figure 2 implies a set of independence conditions:
[…]
As an example for a resulting test case, ID1 states that classification is independent of $T_{E}$ given $H_{s}$, $T_{s}$, and $V_{s}$. A test case could trigger faults at varying $T_{E}$ to verify that the detection probability remains unchanged. As an example for runtime monitoring, ID5 suggests $V_{s}$ and $T_{E}$ should be independent. An additional monitor could track their correlation during operation and trigger an alarm if a threshold is exceeded which would indicate a shift in the assumed operational context (e.g., the probability of a mechanical fault could depend on the environmental temperature which would be a violation of the assumed causal models for this system).
5. Discussion and research agenda
In this vision paper we argue that causal modelling and its mathematical framework have significant potential as an RE activity for ML software system development by systematically integrating prior knowledge into the design. However, based on the experience in our demonstration use case, further research is needed before this vision becomes standard industry practice.
Causal models as complement to natural language requirements. Causal graphs originates from mathematics. We must explore how they can complement current requirements specifications and how they must be adopted for RE. Terms like “treatment”, “confounder”, and “collider” are uncommon in RE and require interpretation.
Criteria for sufficient variable selection. A key challenge is knowing when a causal graphs includes “enough” prior knowledge. We need methods to determine a sufficient set of variables that must be included for a given use case and methods for deciding between competing causal DAGs given a ´´sufficient” set of variables.
Modularisation of ML software systems. Isolating causal mechanisms can guide the modularisation of ML systems, i.e., dividing large monolithic ML models into smaller sub-models.
A common language between different stakeholders of ML software systems. Causal models provide a unified way to communicate prior knowledge and assumptions. Research should explore how this can facilitate coordination between different groups such as data scientists, product experts, and software engineers.
Data requirements derived through causal reasoning. Causal reasoning in RE helps identifying data requirements. Further research should assess to what degree data requirements derived from causal models can enhance ML robustness and reduce data needs compared to traditional RE methods.
Testing and runtime checks. ML software system must align with expected (causal) behaviour. Causal graphs imply independence criteria that lead to testable implications for the runtime behaviour. Research should explore how to translate these into testing strategies and monitors and how reliable such monitors are in practice.
Conclusion
Causal reasoning offers a systematic way to integrate prior knowledge into RE for ML software systems. We outlined a vision and demonstrated a preliminary workflow to derive and argue for low level model and data requirements from high level prior knowledge using causal graphs. We discussed future research activities that are needed to turn this vision into industrial practice.
Acknowledgements.
This project has received funding from the EU’s Horizon research and innovation program under grant agreement No 957197 (vedliot).